OpenStudy (anonymous):

one factor of a polynomial p(x) is (x + 7). what is the numerical value of p(-7) 0 -7 1 7

OpenStudy (anonymous):

So you can write \(p(x)\) as\[p(x)=(x+7) \ q(x)\]now plug -7 in the equation, you get\[p(-7)=(-7+7) \ q(-7)=0 \times \text{a constant number}=0\]